The graph of the equation y =ax^2 + bx + c, where a, b, and c are constants, is a parabola with axis of symmetry x = -3. Find b/a
Answer by richard1234(7193)

The x-coordinate of the vertex is at x = -b/2a, so -3 = -b/2a <==> b/a = 6.